# Non-invertible duality defects in one, two, and three dimensions via gauging spatially modulated symmetry

### v1

@@ 5 Outlook
Stimulated by recent developments of the two types of new symmetries—non-invertible and spatially modulated symmetries, we explore the interplay between them via lattice gauge theories. Introducing two copies of spin models on a lattice defined in one, two, and three dimensions, we demonstrate a way to construct non-invertible duality defects obtained by gauging subsystem or dipole symmetries, which are typical examples of the spatially modulated symmetries that have emerged in a series of studies of fracton topological phases. Such constructions also help one better understand duality defects in higher dimensions in lattice models, which are active areas of research in communities of high energy and condensed matter physics.
@@
In the case of subsystem symmetries, we have constructed non-invertible duality defects whose fusion rules give rise to subsystem charges. In two dimensions, such fusion rules correspond to the subsystem analog of the category theory forming $\text{Rep}(D_{8})$. In three dimensions, fusion rules of the duality defects involve 00-form subsystem charges and higher-form operators, associated with lineon excitations found in the X-cube model. The duality defect is the subsystem analog of $2$-Rep$\left((\mathbb{Z}_{2}^{(1)}\times\mathbb{Z}_{2}^{(1)})\rtimes\mathbb{Z}_{2}^{(0)}\right)$ discussed in the fusion $2$-category theory [51, 39].
@@
In the case of dipole symmetries, especially in two dimensions, a new degree of freedom has been taken into account — dipole algebra, comprised of global and dipole charges. We show that gauging 00-form dipole symmetry with a given dipole algebra in two dimensions gives rise to $1$-form dual dipole algebra where the hierarchical structure of the algebra is inverted. Such a finding is the generalization of the known fact that gauging a global symmetry generates an emergent quantum symmetry, to a dipole symmetry. This new type of duality is also manifest in the fusion rules of the non-invertible duality defects. Indeed, the fusion rule involves 00-form dipole symmetry forming a dipole algebra given a hierarchy and $1$-form dipole symmetry with the hierarchical stricture is inverted (Figs. 5(b) and 6(e)). Our results would comply with recent interests in construction of a unified framework to incorporate various kinds of symmetries, including non-invertible and spatially modulated ones.
One can extend our analysis in many different ways, which we leave for future studies. First, it is interesting to explore gauging more generic group symmetries, in addition to $\mathbb{Z}_{N}$ symmetries. In particular, the non-Abelian dipole symmetry has not been discussed yet. Second, one could study higher multipole symmetries in various dimensions. Especially, in dimensions higher than two, such as three dimensions, it is known that gauging 00-form dipole symmetry with dipole algebra and the one with dual dipole algebra lead to different topological orders compared with the two dimensional analog [53]. Exploring richer structures of (dual) multipole algebras thereof in various dimension provides further insight into the spatially modulated symmetries. Third, systematic investigations on the effects of various boundary conditions should be performed. Specifically, when $N\mid L$ (meaning, when the system size $L$ is not divisible by $N$), we can impose twisted boundary conditions for various symmetries, which affects the ground-state degeneracy of the system; when $N\nmid L$, duality twisted boundary conditions [56] are required, which can lead to exotic topological properties and new insight at critical points [57]. Fourth, it could be an interesting and important problem on its own right to explore phase diagram of the spin models after gauging the spatially modulated symmetries. Especially, after gauging, one could add matter terms coupled with the gauge fields in addition to other terms such as the ones constitute topological model with the modulated symmetries, reminiscent of the Fradkin-Shenker model [58]. Identifying possible phases in the model and investigating the dualities of the modulated symmetries in these phases allow one to gain more fruitful insights on the modulated symmetries. Last but not least, establishing category theory corresponding to the fusion rules of the non-invertible duality defects discovered in this paper will fascinate both physicists and mathematicians.
@@ Acknowledgement
We would like to thank T. Ando, J. H. Han, M. Honda, R. Kalloor, A. Nayak, T. Nakanishi, T. Ohishi, S. Shimamori, K. Shiozaki, and P. Tanay, for helpful discussion. This work is in part supported by KAKENHI-PROJECT-23H01097 and the Koshland Fellowship at the Weizmann Institute of Science.
@@ A.1 Subsystem symmetry in three dimensions
We first give derivation of (57). Recall that an operator $C_{\eta}$ (53) is introduced when constructing duality defects obtained by gauging subsystem symmetries in three dimensions, which has the form
@@
$C_{\eta}$ $=\frac{1}{2^{K}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{xz}_{x}(\hat{z})\right)\left(1+\eta^{zy}_{x}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{xz}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]$
×$\qquad\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{x}(\hat{y})\right)\right]\times\frac{1}{2^{2L_{x}L_{y}L_{z}-K}}\sum_{\gamma_{0}}\eta(\gamma_{0}),$ (141)
@@
$C_{\eta}$ $=\frac{1}{2^{K}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{xz}_{x}(\hat{z})\right)\left(1+\eta^{zy}_{x}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{xz}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]$
×$\qquad\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{x}(\hat{y})\right)\right],$
@@ A.2 Dipole symmetry in two dimensions
We turn to the derivation of (110). When constructing duality defects via gauging dipole symmetry in two dimensions, an operator $C_{1}$ (107) is introduced, which has the form
@@
$C_{1}=\left(\frac{1}{N^{2L_{x}L_{y}-3}}\sum_{\gamma_{0}}\eta(\gamma_{0})\right)\times\left(\frac{1}{N^{3}}\sum_{p,q,r=1}^{N}\sum_{s,t,u=1}^{N}(\zeta_{x,y}^{Z})^{p}(\zeta_{1}^{Z})^{q}(\zeta_{2}^{Z})^{r}(\eta_{x,y}^{Z})^{s}(\eta_{1}^{Z})^{t}(\eta_{2}^{Z})^{u}\right).$ (146)

### v2

@@ 5 Outlook
Stimulated by recent development of the two types of exotic symmetries—noninvertible and spatially modulated symmetries, we explore the interplay between them via lattice gauge theories. Introducing two copies of spin models on a lattice defined in one, two, and three dimensions, we demonstrate a systematic approach to construct noninvertible operators by gauging subsystem or dipole symmetries, which are typical examples of the spatially modulated symmetries that have emerged in fractonic topological phases. Such constructions also help better understand noninvertible operators in higher dimensional lattice models, which are active areas of research in communities of high energy and condensed matter physics.
@@
In the case of subsystem symmetries, we have constructed noninvertible operators whose fusion rules give rise to subsystem charges. In two dimensions, such fusion rules correspond to the subsystem analog of the category theory forming $\text{Rep}(D_{8})$. In three dimensions, fusion rules of the operators involve 00-form subsystem charges and higher-form operators, associated with lineon excitations found in the X-cube model. The operator is the subsystem analog of $2$-Rep$\left((\mathbb{Z}_{2}^{(1)}\times\mathbb{Z}_{2}^{(1)})\rtimes\mathbb{Z}_{2}^{(0)}\right)$ discussed in the fusion $2$-category theory [58, 42].
@@
In the case of dipole symmetries, especially in two dimensions, a new degree of freedom has been taken into account — dipole algebra, comprised of global and dipole charges. We show that gauging 00-form dipole symmetry with a given dipole algebra in two dimensions gives rise to $1$-form dual dipole algebra where the hierarchical structure of the algebra is inverted. Such a finding is the generalization of the known fact that gauging a global symmetry generates an emergent quantum symmetry, to a dipole symmetry. This new type of duality is also manifest in the fusion rules of the noninvertible operators. Indeed, the fusion rule involves 00-form dipole symmetry forming a dipole algebra given a hierarchy and $1$-form dipole symmetry with the hierarchical stricture is inverted (Figs. 5(b) and 6(e)). Our results would comply with recent interests in construction of a unified framework to incorporate various kinds of symmetries, including noninvertible and spatially modulated ones.
One can extend our analysis in many different ways, which we leave for future studies. First, it is interesting to explore gauging more generic group symmetries, in addition to $\mathbb{Z}_{N}$ symmetries. In particular, the non-Abelian dipole symmetry has not been discussed yet.
Second, one could study higher multipole symmetries in various dimensions. Especially, in dimensions higher than two, such as three dimensions, it is known that gauging 00-form dipole symmetry with dipole algebra and the one with dual dipole algebra lead to different topological orders compared with the two dimensional analog [60]. Exploring richer structures of (dual) multipole algebras thereof in various dimension provides further insight into the spatially modulated symmetries.
Third, systematic investigations on the effects of various boundary conditions should be performed. Specifically, when $N\mid L$ (meaning, when the system size $L$ is not divisible by $N$), we can impose twisted boundary conditions for various symmetries, which affects the ground-state degeneracy of the system; when $N\nmid L$, duality twisted boundary conditions [63] are required, which can lead to exotic topological properties and new insight at critical points [64].
Fourth, it could be an interesting and important problem on its own right to explore phase diagram of the spin models after gauging the spatially modulated symmetries. Especially, after gauging, one could add matter terms coupled with the gauge fields in addition to other terms such as the ones constitute topological model with the modulated symmetries, reminiscent of the Fradkin-Shenker model [65]. Identifying possible phases in the model and investigating the dualities of the modulated symmetries in these phases allow one to gain more fruitful insights on the modulated symmetries.
Fifth, it would be intriguing to explore noninvertible duality defects that are generated by gauging the entire symmetry group; by contrast, our noninvertible operators are obtained by gauging a proper subgroup (modulated symmetries for a single copy of spin systems instead of both copies) of the entire symmetry group. Generally, the noninvertible duality defect in $d$ spatial dimension is obtained when the theory is invariant under the gauging $p$-form symmetry. Especially, in the case of the dipole symmetry, the conditions to have such a defect are (i) $d=(p+1)/2$ (ii) The dipole symmetry has the same dipole algebra before and after gauging. It would be also interesting to address whether one can extend our analysis beyond square lattices and examine how dipole symmetry and its algebra is generalized by incorporating other types of spatial symmetries such as rotations and refections.
Last but not least, establishing category theory corresponding to the fusion rules of the noninvertible operators discovered in this paper will fascinate both physicists and mathematicians.
@@ Acknowledgement
We would like to thank T. Ando, J. H. Han, M. Honda, R. Kalloor, A. Nayak, T. Nakanishi, T. Ohishi, T. Saito, S. Shimamori, S. Shimomura, K. Shiozaki, T. Takama, and P. Tanay for helpful discussion. We also thank W. Cao and A. Tiwari for useful comments on the draft and suggesting relevant works that we have missed. This work is in part supported by KAKENHI-PROJECT-23H01097, the Koshland Fellowship at the Weizmann Institute of Science, JST CREST Grant Number JPMJCR24I3.
@@ A.1 Subsystem symmetry in three dimensions
We first give derivation of (63). Recall that an operator $C_{\eta}$ (59) is introduced when constructing operators obtained by gauging subsystem symmetries in three dimensions, which has the form
@@
$C_{\eta}$ $=\frac{1}{2^{K}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{zx}_{x}(\hat{z})\right)\left(1+\eta^{yz}_{y}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{zx}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]$
×$\qquad\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{z}(\hat{y})\right)\right]\times\frac{1}{2^{2L_{x}L_{y}L_{z}-K}}\sum_{\gamma_{0}}\eta(\gamma_{0}),$ (150)
@@
$C_{\eta}$ $=\frac{1}{2^{K}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{zx}_{x}(\hat{z})\right)\left(1+\eta^{yz}_{y}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{zx}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]$
×$\qquad\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{z}(\hat{y})\right)\right],$
@@ A.2 Dipole symmetry in two dimensions
We turn to the derivation of (119). When constructing operators via gauging dipole symmetry in two dimensions, an operator $C_{1}$ (116) is introduced, which has the form
@@
$C_{1}=\left(\frac{1}{N^{2L_{x}L_{y}-3}}\sum_{\gamma_{0}}\eta(\gamma_{0})\right)\times\left(\frac{1}{N^{3}}\sum_{p,q,r=1}^{N}\sum_{s,t,u=1}^{N}(\xi_{x,y}^{Z})^{p}(\xi_{1}^{Z})^{q}(\xi_{2}^{Z})^{r}(\eta_{x,y}^{Z})^{s}(\eta_{1}^{Z})^{t}(\eta_{2}^{Z})^{u}\right).$ (155)